OpenStudy (anonymous):

Any body good at solving 1/2 angle questions thetha=480 Find: sin1/2thetha cos1/2thetha tan1/2thetha

OpenStudy (anonymous):

\[\theta = 480 \] \[\sin \frac{ \theta }{ 2 }= \sin \frac{ 480 }{ 2} = \sin 240\]

OpenStudy (anonymous):

cos 240 = tan 240 =

OpenStudy (anonymous):

uhm so the answer would be for sin -rad3/2 cos1/2 =-1/2 tan1/2= rad 3

OpenStudy (anonymous):

sin 240deg = -0.86602540378